# Learning Prompt with Distribution-Based Feature Replay for Few-Shot Class-Incremental Learning

### v1

@@ 10 Analysis Various Shot Numbers
In the main paper, we have compared the performance of our LP-DiF with SOTA methods and the Joint-LP (upper bound) under $5$-shot for each incremental class. To further demonstrate the superiority of our approach, we conducted experiments under various shot numbers of incremental classes. Fig. 8 show the comparison results with BiDistFSCIL [56] and Joint-LP on CUB-200 under (a) $2$-shot, (b) $5$-shot (default in main paper), (c) $10$-shot and (d) $15$-shot. Obviously, across all the shot number settings, our LP-DiF consistently outperforms BiDistFSCIL significantly, and its performance is very close to the upper bound. This result demonstrates that, regardless of the shot numbers of incremental classes, our LP-DiF presents satisfactory performance and the ability to resist catastrophic forgetting.

### v3

@@ IV-D Ablation Studies and Analysis
Analysis on Shot Numbers. To further demonstrate the superiority of our approach, we conducted experiments under various shot numbers of incremental classes. Fig. 8 show the comparison results with BiDistFSCIL [60] and Joint-LP on CUB-200 under (a) $2$-shot, (b) $5$-shot, (c) $10$-shot and (d) $15$-shot. Obviously, across all the shot number settings, our LP-DiF consistently outperforms BiDistFSCIL significantly, and its performance is very close to the upper bound. This result demonstrates that, regardless of the shot numbers of incremental classes, our LP-DiF presents satisfactory performance and the ability to resist catastrophic forgetting.
Analysis on Effect of VAE. From the results presented in Tab. VIII, one can observe that using both synthesize features (SF) and real features (RF) for estimating the Gaussian distribution (GD), as compared to using only real features, can achieve higher performance. To elucidate the quality of the features synthesized by the VAE, we conducted the following analysis on mini-ImageNet: we calculate the difference between the estimated GD of each incremental class and corresponding reference GD (ideal GD, i.e., computed using the full training images of that class) by KL divergence $D_{\text{KL}}$
$D_{\text{KL}}(P|Q)=\frac{1}{2}\sum_{i=1}^{n}\left(\frac{\sigma_{p_{i}}^{2}}{\sigma_{q_{i}}^{2}}+\frac{(\mu_{q_{i}}-\mu_{p_{i}})^{2}}{\sigma_{q_{i}}^{2}}-1+\ln\left(\frac{\sigma_{q_{i}}^{2}}{\sigma_{p_{i}}^{2}}\right)\right),$ (12)
where $P$ and $Q$ represent the ideal GD and the estimated GD respectively, and $n$ represents the dimension of GD. Statistically, lower $D_{\text{KL}}$ indicates that the estimated GD is closer to reference GD. Fig 7 shows the results on mini-ImageNet. Note that for most classes, the GD estimated using both SF and RF is closer to the reference distribution, indicating that SF can enrich more class-relevant information.
Training Time and Model Size. Compared to existing ResNet-based FSCIL methods, our LP-DiF is based on the heavier model (i.e., CLIP), which may raise concerns about model training efficiency and memory overhead. However, since LP-DiF only trains lightweight prompt vectors and a few layers of MLP in the VAE, it does not incur excessive computational costs. Here, we offer some quantitative results for reference: For the training time, with 8×8$\times$8 × 2080ti GPUs, for each incremental session, training LP-DiF takes about 4.5 minutes (1.8 minutes for training VAE and 2.7 minutes for training prompts). In comparison, BiDistFSCIL [5] takes about 3.3 minutes for training for the same epochs. For the volume of trainable parameters, existed FSCIL methods relied on ResNet require training about 11.3M parameters for ResNet-18, respectively. However, LP-DiF only needs to train about 7.4M parameters for prompts and MLPs. Thus, our LP-DiF achieved significant performance gain with acceptable addition on training time and lower volume of trainable parameters.
V Conclusion
In this paper, we studied the FSCIL problem by introducing V-L pretrained model, and proposed Learning Prompt with Distribution-based Feature replay (LP-DiF). Specifically, prompt tuning is involved to adaptively capture the knowledge of each session. To alleviate catastrophic forgetting, we established a feature-level distribution for each class, which is estimated by both real features of training images and synthesized features generated by a VAE decoder. Then, pseudo features are sampled from old-class distributions, and combined with the training set of current session to train the prompts jointly. Extensive experiments show that our LP-DiF achieves the new state-of-the-art in the FSCIL task.